k^2 - 25 = 0
please help me solve useing the factor of 0 please show step by step


i have k=5 but i do it off my head and dont know how to write it down in the long version

Answer by mananth(16949)

k^2-25=0
this is of the form a^2-b^2= (a+b)(a-b)
(k+5)(k-5)=0
k= + / - 5